OpenStudy (anonymous):

What is the equation (in slope-intercept form) of a line that passes through the point (3, -2) and has a slope of -2?

OpenStudy (anonymous):

Choices: y = 2x + 4 y = -2x + 4 y + 2 = -2(x + 3) 2x + y + 4 = 0

OpenStudy (anonymous):

Start with the point-slope form of the equation. m is the slope and (x_1, y_1) is the point. Plug in the numbers \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

Then distribute "m" and add y_1 to get into slope intercept form. \[y=mx+b\]

OpenStudy (anonymous):

ok thanks! I get it now